OpenStudy (anonymous):

The ratio of two numbers is 5 to 3. The sum of the numbers is 16. What are the numbers?

OpenStudy (aravindg):

a/b=5/3 a+b=16 Substitute for a or b from first equation to second equation.

OpenStudy (anonymous):

10 and 6 ? @AravindG

OpenStudy (aravindg):

Yep :)

OpenStudy (anonymous):

Thank you

OpenStudy (aravindg):

yw :)